coordinate geometry: find two points on the perpendicular bisector of AB.
verify your answer results by showing each point is equidistant from A and B
A(0,0),B(0,4)

Answer by Alan3354(69443)

Since both points are on the y-axis, the perpendicular bisector will be parallel to the x-axis and 2 units from it.
Any point on y = 2 fits.
(1,2)
(5,2)
(-55,2)
(x,2) works.
-----------
To find the distance for 2 of the points:
(1,2) from (0,0)
s^2 = diffy^2 + diffx^2 = 1 + 4
s^2 = 5
(1,2) from (0,4)
s^2 = 1 + 4
Same distance (if the squares are equal, they're equal)
---------
---------
(-55,2) from (0,0)
s^2 = 55^2 + 4 = 3029
(-55,2) from (0,4)
s^2 = 3025 + 4 = 3029 --> same distance sqrt(3029)